OpenStudy (anonymous):

what polynomial equals (x+6)(2x-3) how do you know please explain

OpenStudy (anonymous):

You multiply x * 2x = ? x * -3 = ? 6 * 2x = ? 6 * -3 = ? and combine like terms. Also known as FOIL First Outside Inside Last

OpenStudy (anonymous):

2x^2 + 12X +3

OpenStudy (anonymous):

Can you show your work so I can better explain where you went wrong

OpenStudy (anonymous):

x * 2x = 2x^2 x * -3 = ? 6 * 2x = 12x 6 * -3 =?

OpenStudy (anonymous):

3 =Xx-3 -18=6x-3

OpenStudy (anonymous):

close a negative and a positive = a negative so x * -3 = -3x

OpenStudy (anonymous):

Also you were forgetting to keep the variable x when multiplying x and -3 Now we know that x * 2x = 2x^2 x * -3 = -3x 6 * 2x = 12x 6 * -3 = -18 we can combine like terms

OpenStudy (anonymous):

Do you see any like terms?